OpenStudy (anonymous):

A pair of equations is shown below: y = 8x − 9 y = 4x − 1 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

i will fan and medal

OpenStudy (michele_laino):

hint: you can solve using the elimination method

OpenStudy (anonymous):

i put it in subtitution method but this is all i got: part A: im am using the substitution method. both of these equations are solved for y. im going to substitute the first equation into the second equation. 8x-9=4x-1. im going to add 8x to both sides. 8x+8x-9=4x+8x-1. 16x-9=12x-1 add 9 to both sides. 16x=12. divide both sides by 16. 16/12=

OpenStudy (anonymous):

i dont know if im doing it right though

OpenStudy (michele_laino):

here is the first step: \[\Large 8x - 9 = y = 4x - 1\]

OpenStudy (anonymous):

ok. know i add 8 to both sides right?

OpenStudy (michele_laino):

we can add 9 to both sides, so we get: \[\Large 8x = 4x - 1 + 9\]

OpenStudy (anonymous):

ok. 8x=4x-10 right?

OpenStudy (michele_laino):

8x=4x+8

OpenStudy (anonymous):

wait so the one is a negative? i have been getting confused with that. so what would be the next step?

OpenStudy (michele_laino):

next step: we can subtract 4x at both sides, so we get: \[\Large \begin{gathered} 8x - 4x = 8 \hfill \\ 4x = 8 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok. then we divide?

OpenStudy (michele_laino):

yes! we can divide both sides by 4

OpenStudy (michele_laino):